OpenStudy (anonymous):

If \(\quad \large abc = 1\), \(\quad\) then find: \[\large \color{green}{(\frac{1}{1+a + b^{-1}} + \frac{1}{1 + b + c^{-1}} + \frac{1}{1 + c + a^{-1}})}\] \[\large \color{blue}{\mathit{Answer : 1}}\]

OpenStudy (anonymous):

so far i got.. \[\large \frac{6+3a+3b+3c+3bc+3ac+3ab+a^2b+b^2c+c^2a}{(b+ab+a)(c+bc+1)(a+ac+1)}\] lol

OpenStudy (anonymous):

Ha ha ha... Try hard..

OpenStudy (anonymous):

=( what do you suggest to do next?

OpenStudy (anonymous):

Going totally wrong.. Even Einstein will fail to solve this what you are doing...

OpenStudy (anonymous):

all denominators are equal,,thats what i got so far..

OpenStudy (anonymous):

How you made all the denominator equal??

OpenStudy (anonymous):

ohh sorry not equal at the moment,,we have this on simplification:

OpenStudy (anonymous):

(a+b+c)/(1+(a-1) +c)

OpenStudy (anonymous):

\[ \frac{1}{1+a+b^{-1}}\cdot\frac {bc} {bc}=\frac{bc}{bc+abc+c}=\frac{a^{-1}}{1+c+a^{-1}} \]

OpenStudy (anonymous):

no am sorry,,am getting confused..gimme 1 min..

OpenStudy (anonymous):

After that...??

OpenStudy (anonymous):

just let a=x/y b=y/z c=z/x

OpenStudy (anonymous):

Yes go ahead @mukushla ..

OpenStudy (anonymous):

then u have \[\frac{y}{x+y+z}+\frac{z}{x+y+z}+\frac{x}{x+y+z}=\frac{x+y+z}{x+y+z}=1\]

OpenStudy (anonymous):

got it..1 min i'll post

OpenStudy (anonymous):

I know one method to solve it.. Now I know two.. Thank you very much @mukushla

OpenStudy (anonymous):

1/(1+a+(b-1)) = b/(b+(c-1)+1) 1/(1+c+(a-1) = a/(a+(b-1)+1) = ab/(b+(c-1)+1) = (c-1)/(b+(c-1)+1) adding all easily =1

OpenStudy (accessdenied):

If you go with @nphuongsun93 's attempt and expand those nasty multiplications, it does work out to be the same on top and bottom, = 1 It's just really hard to stay organized / avoid mistakes, and it takes a lot longer than the above methods. D:

OpenStudy (anonymous):

I have another solution.. Everyone should know that: Leave the first term as such, multiply and divide by \(b^{-1}\) to numerator and denominator of middle term and for third term, by taking LCD just solved the denominator, You will get: \[= \frac{1}{1 + a + b^{-1}} + \frac{b^{-1}}{b^{-1} + 1 + b^{-1}c^{-1}} + \frac{a}{a + ac + 1}\] Using abc = 1 \[= \frac{1}{1 + a + b^{-1}} + \frac{b^{-1}}{1 + b^{-1} + a} + \frac{a}{a + b^{-1} + 1}\] \[= \frac{1 + a + b^{-1}}{1 + a + b^{-1}} \implies = 1\]

OpenStudy (anonymous):

hardest way for the hardworkers!!